OpenStudy (anonymous):

Before taking a typing course, Frank could type 27 words per minute. By the end of the course, he was able to type 77 words per minute. Find the percent increase.

OpenStudy (anonymous):

Solve. Round to the nearest tenth, if necessary.

OpenStudy (anonymous):

50/27 * 100

OpenStudy (amistre64):

77-27 ------ * 100 27 indiana got it gooood

OpenStudy (anonymous):

A. 185.2% B. 74% C. 35.1% D. 64.9%

OpenStudy (anonymous):

None of you guys answers are on here :(

OpenStudy (anonymous):

A. 185.2% (77 - 27) / 27 = 50/ 27 = 185.2 %

OpenStudy (amistre64):

it helps if you can do some subtracting and multiplying to get yourself to the end product :/ instead of relying upon the masses to do it for you

OpenStudy (anonymous):

I wish I can give you more medals for that.